A toddler takes off running down the sidewalk at 210 ​ft/min. One minute ​later, a worried mother runs after the child at 470 ​ft/min. After how many minutes will the mother overtake the​ toddler?
Answer by ankor@dixie-net.com(22740)

A toddler takes off running down the sidewalk at 210 ​ft/min.
One minute ​later, a worried mother runs after the child at 470 ​ft/min.
After how many minutes will the mother overtake the​ toddler?
:
Let t = running time of the mother
The toddler has a 1 min head start, therefore
(t+1) = running time of the toddler
:
When Mom catches child, they will have run the same distance
write a dist equation; dist = speed * time
Mom dist = child dist
470t = 210(t+1)
470t = 210t + 210
470t - 210t = 210
260t = 210
t = 210/260
t = .81 min or .81(60) = 48.6 sec, that's a fast Mommy!